The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 33 June 13, 1999

Contents

News  1
Abstracts of 7 refereed papers  3
Conference contribution abstract 6
Thesis abstract  7

News

In this issue we would like to commemorate the recent passing of two prominent Magellanic Clouds researchers, Dr. Barry Lasker (STScI, Baltimore, USA) and Dr. Rebecca Elson, (IoA, Cambridge, UK).

Barry Lasker (1939–1999)

Barry Lasker received his Bachelor of Science degree in physics from Yale in 1961, Master’s degree in 1963, and a Ph.D. in 1964 in astrophysics from Princeton. Barry has been a postdoc at the Hale Observatories, a professor of astronomy at the University of Michigan, a staff astronomer at CTIO, and a staff scientist at the STScI. A detailed description of Barry’s career can be found in the cover article of the STScI Newsletter, June 1999 issue. Barry’s most important contribution to astronomy is undoubtedly his work on the Guide Star Catalog and the Digital Sky Survey, for which he has been awarded the 1999 Van Biesbroeck Prize from the American Astronomical Society and the 1999 Muhlmann Award from the Astronomical Society of the Pacific.

What has escaped people’s attention is Barry’s significant contribution to the study of the Large Magellanic Cloud, which was his main research topic at CTIO shortly before he left for the STScI. He made the first photographic survey of the LMC in multiple emission lines – Hα, [O III], and [S II], using the Curtis Schmidt telescope (1979, CTIO Contribution No. 127). The survey was an invaluable source of information for studies of ionized gas in the LMC. Barry has studied WR ring nebulae, supernova remnants, and superbubbles in the LMC using multiple nebular line morphologies and kinematics. Even after Barry moved to the STScI, he still remembered the Magellanic Clouds, and he was a leading US member in the the HST program “Parallel high resolution imaging of diffuse objects in the Magellanic Clouds”.

Barry’s pioneer work on shell nebulae in the LMC has inspired many researchers in this field. Barry will be sorely missed by the Magellanic Cloud researchers, especially those of us who are privileged to have worked with Barry and shared his 4m prime focus images of LMC superbubbles.

Rebecca Elson (1960–1999)

Becky Elson received her M.Sc. from the University of British Columbia, Vancouver, Canada (1982), and her Ph.D. from the Institute of Astronomy, Cambridge, UK (1986). After postdoctoral fellowships at the Institute of Advanced Study, Princeton, USA (1986–1989), at the Bunting Institute, Radcliffe College, USA, and the Harvard-Smithsonian Center for Astrophysics, Cambridge, USA, she returned to the Institute of Astronomy, Cambridge, UK, where she worked from 1991 to 1999.^{1} During her distinguished research career she published 34 refereed papers, an ARA&A review on the dynamical evolution of star clusters, and numerous conference reviews and contributions. Becky’s research concentrated on stellar populations and star clusters in different environments. She worked on globular cluster systems of spirals and ellipticals, on Milky Way globular clusters, and on rich star clusters and field stars in the Magellanic Clouds.

Highlights of Becky’s work on the Magellanic Clouds include extensive studies of the structural parameters of Magellanic Cloud clusters, an SWB-type age calibration of integrated UBV colors of clusters in LMC and SMC (in part with Mike Fall and Ken Freeman), which remain milestones in the field. Her recent work concentrated on deep luminosity functions, binary fractions, and mass segregation from HST data (with Gerry Gilmore and others). Among the highlights of her research unrelated to the Magellanic Clouds are the first determination of structural parameters for globular clusters outside the Local Group (with David Schade), and a study of the most distant globular cluster system ever observed around an elliptical galaxy in the Hubble Deep Field (with Gerry Gilmore and Basilio Santiago).

To honor Becky’s important contributions to star cluster research we dedicated the one-day topical session on “Extragalactic Star Clusters” at the Centennial AAS meeting two weeks ago in Chicago to her. The proceedings of the X. Canary Islands Winter School on “Globular Clusters”, where she lectured on “Stellar Dynamics in Globular Clusters”, will also be dedicated to her memory. Becky will be missed by her friends and colleagues.

Eva Grebel & You-Hua Chu

1We thank Gerry Gilmore for kindly contributing Rebecca Elson’s vita.

Abstracts of Refereed Papers

Superbubble evolution including the star-forming clouds: Is it possible to reconcile LMC observations with model predictions?

S. Silich^{1} and J. Franco^{2}

1 Main Astronomical Observatory National Academy of Sciences of Ukraine, 252650 Kyiv, Golosiiv, Ukraine 2 Instituto de Astronomía-UNAM, Apdo. Postal 70-264, 04510 México D. F., Mexico

Here we present a possible solution to the apparent discrepancy between the observed properties of LMC bubbles and the standard, constant density bubble model. A two-dimensional model of a wind-driven bubble expanding from a flattened giant molecular cloud is examined. We conclude that the expansion velocities derived from spherically symmetric models are not always applicable to elongated young bubbles seen almost face-on due to the LMC orientation. In addition, an observational test to differentiate between spherical and elongated bubbles seen face-on is discussed.

Accepted by: Astrophysical Journal (September 1999, Vol. 522) For preprints, contact silich@mao.kiev.ua

Also available from the URL http://xxx.lanl.gov/astro-ph/9905167

ISO SWS/LWS observations of SN 1987A

P. Lundqvist^{1}, J. Sollerman^{1}, C. Kozma^{1}, B. Larsson^{1}, J. Spyromilio^{2} A.P.S. Crotts^{3}, J. Danziger^{4} and D. Kunze^{5}

1 Stockholm Observatory, SE-133 36 Saltsjöbaden, Sweden 2 European Southern Observatory, Karl-Schwarzschild-Strasse 2, D-85748 Garching, Germany 3 Columbia University, Dept. of Astronomy, 538 W. 120th Street, New York, USA 4 Osservatorio Astronomico, Via G.B. Tiepolo 11, I-34131 Trieste, Italy 5 Max-Planck-Institut für extraterrestrische Physik, Postfach 1603, D-85740 Garching, Germany

We report on observations of SN 1987A with ISO SWS/LWS made 9−11 years after the explosion. No emission from the supernova was seen. In particular, the upper limits on the fluxes of [Fe I] 24.05μm and [Fe II] 25.99μm on day 3 999 are ∼ 1.1 Jy and ∼ 1.4 Jy, respectively. Assuming a homogeneous distribution of ^{44}Ti inside 2 000 km s^{−1}, we have made theoretical models to estimate the mass of ejected^{44}Ti. Assessing various uncertainties of the model, we obtain an upper limit of ≃ 1.5×10^{−4} M_{⊙}. The implications of this are discussed.

The LWS data display continuum emission as well as nebular lines of [O I], [C II] and [O III] from neighboring photoexcited regions in the LMC. The [O III] lines indicate an electron density of 120 ± 75 cm^{−3}, and the continuum can be explained by dust with a temperature of ∼ 37 K. A second dust component with ∼ 10 K may also be present.

Accepted by: Astronomy & Astrophysics For preprints, contact peter@astro.su.se

Also available from the URL ftp://www.astro.su.se/pub/supernova/letal99 87A/

Infrared Photometry of Red Supergiants in Young Clusters in the Magellanic Clouds

Stefan C. Keller^{1}

1 Research School of Astronomy and Astrophysics, The Australian National University, Weston Creek P.O., ACT 2611, Australia.

We present broad-band infrared photometry for 52 late-type supergiants in the young Magellanic Clouds clusters NGC 330, NGC 1818, NGC 2004 and NGC 2100. Standard models are seen to differ in the temperature they predict for the red supergiant population on the order of 300K. It appears that these differences most probably due to the calibration of the mixing-length parameter, α_{P}, in the outermost layers of the stellar envelope. Due to the apparent model dependent nature of α_{P} we do not quantitatively compare α_{P} between models. Qualitatively, we find that α_{P} decreases with increased stellar mass within standard models. We do not find evidence for a metallicity dependence of α_{P}.

Accepted by: The Astronomical Journal (astro-ph 9905030)

For preprints, contact stefan@mso.anu.edu.au
Also available from the URL http://msowww.anu.edu.au/~stefan/work.html

Theoretical Models for Classical Cepheids: IV. Mean Magnitudes and Colors and the Evaluation of Distance, Reddening and Metallicity.

F. Caputo, M. Marconi, V. Ripepi

Astron. Obs. of Capodimonte, Italy

We discuss the metallicity effect on the theoretical visual and near-infrared PL and PLC relations of classical Cepheids, as based on nonlinear, nonlocal and time–dependent convective pulsating models at varying chemical composition. In view of the two usual methods of averaging (magnitude-weighted and intensity-weighted) observed magnitudes and colors over the full pulsation cycle, we briefly discuss the differences between static and mean quantities. We show that the behavior of the synthetic mean magnitudes and colors fully reproduces the observed trend of Galactic Cepheids, supporting the validity of the model predictions. In the second part of the paper we show how the estimate of the mean reddening and true distance modulus of a galaxy from Cepheid VK photometry depend on the adopted metal content, in the sense that larger metallicities drive the host galaxy to lower extinctions and distances. Conversely, self-consistent estimates of the Cepheid mean reddening, distance and metallicity may be derived if three-filter data are taken into account. By applying the theoretical PL and PLC relations to available BVK data of Cepheids in the Magellanic Clouds we eventually obtain Z ∼ 0.008, E(B-V) ∼ 0.02 mag, DM ∼ 18.63 mag for LMC and Z ∼ 0.004, E(B-V) ∼ 0.01 mag., DM ∼ 19.16 mag. for SMC. The discrepancy between such reddenings and the current values based on BVI data is briefly discussed.

Accepted by: The Astrophysical Journal

For preprints, contact marcella@na.astro.it

Spectroscopic analysis of the candidate globular clusters NGC 1928 and 1939 in the Large Magellanic Cloud

C.M. Dutra^{1}, E. Bica^{1}, J.J. Claria^{2}, A.E. Piatti^{2}

1 Departamento de Astronomia, Instituto de Fisica, UFRGS, C.P. 15051, 91501-970 Porto Alegre RS, Brazil 2 Observatorio Astronomico de Cordoba, Laprida 854, 5000, Cordoba, Argentina.

The integrated spectral properties in the range 3600-6700 Å of the candidate old clusters NGC 1928 and 1939 in the LMC bar are compared with those of old- and intermediate-age reference LMC clusters, the properties of which are better established. It has been possible to infer the age of the sample clusters by means of absorption features and the continuum distribution, in particular in the plane W_{M} × W_{B} (where W_{B} is the average of Hδ, Hγ, and Hβ equivalent widths, and W_{M} that of Ca II K, G band and Mg i). The results indicate that NGC 1928 and 1939 are compatible with old clusters. The metallicity is derived with respect to galactic globular cluster templates: [Fe/H] ∼ −1.2 and −2.0 for NGC 1928 and 1939, respectively. We also discuss the census of Population II clusters in the LMC, their spatial distribution and the possibility of a LMC core and a transient morphological classification for interacting late-type disc galaxies.

Published: MNRAS, 305, 373, 1999 For preprints, contact dutra@if.ufrgs.br

A New Kinematic Distance Estimator to the LMC

A. Gould^{1}

1 Ohio State University, Department of Astronomy, Columbus, OH 43210, USA

The distance to the Large Magellanic Cloud (LMC) can be directly determined by measuring three of its properties, its radial-velocity field, its mean proper motion, and the position angle ϕ_{ph} of its photometric line of nodes. Statistical errors of ∼ 2% are feasible based on proper motions obtained with any of several proposed astrometry satellites, the first possibility being the Full-Sky Astrometric Mapping Explorer (FAME). The largest source of systematic error is likely to be in the determination of ϕ_{ph}. I suggest two independent methods to measure ϕ_{ph}, one based on counts of clump giants and the other on photometry of clump giants. I briefly discuss a variety of methods to test for other sources of systematic errors.

Submitted to: The Astrophysical Journal

For preprints, contact Andy Gould, gould@astronomy.ohio-state.edu

Effect of the Milky Way on Magellanic Cloud structure

Martin D. Weinberg^{1}

1 Department of Physics & Astronomy, University of Massachusetts, Amherst, MA 01003-4525, USA

A combination of analytic models and n-body simulations implies that the structural evolution of the Large Magellanic Cloud (LMC) is dominated by its dynamical interaction with the Milky Way.

Although expected at some level, the scope of the involvement has significant observational consequences. First, LMC disk orbits are torqued out of the disk plane, thickening the disk and populating a spheroid. The torque results from direct forcing by the Milky Way tide and, indirectly, from the drag between the LMC disk and its halo resulting from the induced precession of the LMC disk. The latter is a newly reported mechanism that can affect all satellite interactions. However, the overall torque can not isotropize the stellar orbits and their kinematics remains disk-like. Such a kinematic signature is observed for nearly all LMC populations. The extended disk distribution is predicted to increase the microlensing toward the LMC. Second, the disk’s binding energy slowly decreases during this process, puffing up and priming the outer regions for subsequent tidal stripping. Because the tidally stripped debris will be spatially extended, the distribution of stripped stars is much more extended than the HI Magellanic Stream. This is consistent with upper limits to stellar densities in the gas stream and suggests a different strategy for detecting the stripped stars. And, finally, the mass loss over several LMC orbits is predicted by n-body simulation and the debris extends to tens of kiloparsecs from the tidal boundary. Although the overall space density of the stripped stars is low, possible existence of such intervening populations have been recently reported and may be detectable using 2MASS.

Submitted to: The Astrophysical Journal

For preprints, contact weinberg@phast.umass.edu
Also available from the URL http://www-astro.phast.umass.edu/~weinberg/weinberg-pubs.html

Abstracts of Non-Refereed Papers

Morphology and Evolution of Galactic and Magellanic Cloud Planetary Nebulae

L. Stanghellini^{1,2}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore MD 21218, USA 2 Affiliated to the Astrophysics Division, Space Science Department of ESA

Planetary nebulae (PNe) exist in a range of different morphologies, from very simple and symmetric round shells, to elliptical, bipolar, and even quadrupolar shapes. They present extremely complex ensembles of filaments, knots, ansae, and shell multiplicity. It is then overwhelmingly complicated to derive reasonable evolutionary paths to justify the observed shapes of PNe. The confrontation between the evolution of the shells and that of the central stars is needed to understand the origin of the morphological variety. We present some background and recent results on the correlations between PN morphology and PN nuclei (PNNi) evolution, including a study on the Magellanic Cloud PNe.

To appear in: Astrophysical Dynamics, Conference to commemorate the work of Franz D. Kahn. To be published in Astrophysics and Space Science (Kluwer). For preprints, contact lstanghe@stsci.edu

Thesis Abstract

Reconstruction of Star Formation Histories of Resolved Stellar Populations

Andrew E. Dolphin^{1}

1 Department of Astronomy, University of Washington, Box 351580, Seattle, WA 98185-1580, USA

A method for the study of star formation histories of resolved galaxies is presented. It is tested with both simulated data of a composite population and observed data of a globular cluster, which is expected to be a single population. The method is then used to study the star formation histories of WLM, the LMC, and LGS 3.

For WLM, HST F555W and F814W photometry of a portion of the galaxy is presented, with the chips aligned such that sample of both the bar and halo populations were present in the data. The distance modulus for the globular cluster is calculated to be 24.72 ± 0.09 from color-magnitude diagram fitting, with [Fe/H] of -1.51 ± 0.09 and age of 14.8 ± 0.7 Gyr. The distance is consistent with distances calculated for the cluster HB and the field star TRGB. Large-scale star formation in the galaxy began about 12 Gyr with a short but large burst. Intermediate age star formation rates are uncertain, but the bar has shown increased star formation during the past 1.5 Gyr.

The LMC data is ground-based UBV photometry of two fields in the northern part of the disk. A distance modulus of 18.40 ± 0.05 is calculated. Large-scale star formation began about 12 Gyr ago in this part of the LMC with a large burst, and slowly decreased. A recent burst began between 1.0 and 2.5 Gyr ago, and shows different recent star formation rates in the two fields.

The LGS 3 data is HST-based F555W and F814W photometry, and includes most of the galaxy’s core and part of the galaxy’s halo. The distance modulus is calculated to be 23.89 ± 0.05, and is consistent with HB and TRGB distances. LGS 3 had two bursts of star formation, one beginning 14.5 ± 0.8 Gyr ago and one beginning between 2.5 and 5 Gyr ago, and ending at least 200 Myr ago. The recent burst was constrained to the body of the galaxy.

Ph.D. Thesis, University of Washington, June 1999 Advisor: Paul Hodge

For more information, contact dolphin@astro.washington.edu